# A Symmetric Regressor for MRI-Based Assessment of Striatal Dopamine Transporter Uptake in Parkinson’s Disease

### v1

@@ 4.3 SBR Prediction Performance
The proposed symmetric regressor resulted in an average correlation of $0.7291$, which is a considerable increase from the standard regressor’s correlation (i.e., $0.6674$). Note that, this resultant correlation is within the acceptable range in the medical field. The resultant MAE is also improved from $0.7675$ to $0.6741$ by the proposed method. We compared the results of $VGG$-small, $Symmetric$ $VGG$, and $Symmetric$ $VGG$ + $L_{\text{sym}}$ using Student’s $t$-test, which revealed that the improvements by $Symmetric$ $VGG$ (w.r.t $VGG$-$small$) and $Symmetric$ $VGG+L_{\text{sym}}$ (w.r.t $Symmetric$ $VGG$) were statistically significant ($p$ value $<0.01$ and $p$ value $<0.03$, respectively).
We also visually investigated the extracted features by the proposed model to evaluate the discriminative nature of the learned representation of our data. We do so by mapping the extracted features into 2D space using t-SNE [32]. We plot the resultant 2D features in Fig. 4. Each point in this plot is color-graded based on its true SBR value. We also plot the features extracted by the standard regressor for comparison. It is clear from the plots that the proposed model could learn comparatively more powerful feature representation that is useful for discriminating among different SBR ranges.
Fig. 5 presents one example for each of the high and low SBR cases. For the low SBR case, gradients of the low SBR class were backpropagated to obtain the activation maps (Fig. 5-top). On the other hand, gradients of the high SBR class were used for the high SBR case (Fig. 5-bottom). Generally, the activation maps for high SBR prediction highlighted the nigrosome, parts of the substantia nigra, and parts of the cerebral peduncle. Contrarily, the activation maps for low SBR mainly highlighted the red nucleus region with a small minor focus on parts of the substantia nigra. The low SBR activation maps usually did not show any attention near the nigrosome because the nigrosome hyperintensity is lost in the low SBR cases. This coincides with previous medical research [14, 18] where nigrosome hyperintensity was used as an identifier for normal/abnormal SPECT findings.
@@
4.4 Uncertainty Prediction Performance
For plotting the curve, we obtain the (sharpness, CP) points by varying the interval from $0.\sigma$ to $m\sigma$, given the original interval $\sigma$ from MC dropout. Here, $m>0$ is defined such that the CP reaches $100\%$. Fig. 6 presents the resultant sharpness-CP curves. The figure presents three curves given by the MC intervals of the standard VGG regressor, the MC intervals of the proposed regressor, and our symmetric MC intervals of the proposed regressor. Both intervals of the proposed regressor have a greater AUC than the standard regressor.
The interval we proposed using symmetric MC dropout gave a slightly better AUC than the usual MC interval. For the final regressor, the curves of the symmetric MC and MC intervals are almost similar except only in a few places. However, the proposed MC is still advantageous because high CP positions are usually required for practical application in medical sites. In such high CP positions, the proposed MC interval gave improved sharpness overall. To highlight this advantage, we present the optimal (sharpness, CP) for different prediction intervals in Table 4. In this table, we also include the sharpness at $95\%$ CP, which is the most commonly used confidence rate. At this high CP, the symmetric MC interval gave a sharpness of $74.9\%$ while the typical MC interval gave $71.9\%$.

### v2

@@ 5.3 Explainability Analysis
5.4 Uncertainty Analysis
Based on the sharpness-CP curve in Fig 4, we comparatively analyze the uncertainty of the regressor models in Table 5. We present our discussion on this in two steps. First, we compare the uncertainty between the standard regressor and the proposed symmetric regressor. Then, we present the advantage of the symmetric MC uncertainty estimation method over the typical MC method. We base our comparison on the sharpness (i.e. precision) and coverage (i.e. recall) measures as mentioned earlier in Section 4.4. Besides presenting the optimal sharpness and coverage, we also compare the sharpness at a high coverage (see Table 5). Specifically, we selected the point of $95\%$ coverage, which is the most commonly used confidence rate.
From Table 5, we can see that the prediction intervals of the proposed regressor had higher optimal sharpness and coverage compared with the standard regressor. The average normalized width of the standard regressor’s prediction interval was $0.191$ with a $81.6\%$ probability of covering the true SBR within it. On the other hand, the symmetric regressor’s interval width was $0.187$ with a $84.3\%$ coverage probability. The higher certainty and precision of the proposed regressor was also prominent at the $95\%$ coverage.
The interval estimated by the symmetric MC dropout method had a an optimal width of $0.187$ with a coverage of $84.9\%$, which is only slightly improved from the typical MC method. However, the improved and sharper estimation of the interval is prominent at higher coverage. For example, at $95\%$ coverage, the symmetric MC method gave an average interval of width $0.251$, which is a notable improvement from the standard MC interval width of $0.281$. Therefore, the symmetric regressor coupled with the symmetric MC uncertainty can be used in the clinical sites more reliably.
@@ 6 Conclusion
Acknowledgement
Funding: This work was supported in part by Hankuk University of Foreign Studies (HUFS) Basic Research Fund, South Korea (No. 20231140001) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1F1A1064530) and grant No. 02–2023-0011 from the Seoul National University Bundang Hospital Research Fund (corresponding authors: Il Dong Yun and Yun Jung Bae).